# Nonreciprocal tripartite entanglement and asymmetric Einstein-Podolsky-Rosen steering via directional quantum squeezing

### v1

@@ II Theoretical model and system dynamics
For the CW input case, there is merely a coherent driving field with frequency $\omega_{d}/2$ applied to the waveguide. In this situation, the Hamiltonian of the whole system reads
$\hat{H}_{\circlearrowright}=$ $\hat{H}_{\textrm{sys}}+\hat{H}_{\textrm{dr},\circlearrowright},$
$\hat{H}_{\textrm{sys}}=$ $\Delta_{c}\hat{a}_{\circlearrowright}^{\dagger}\hat{a}_{\circlearrowright}+\sum_{j=1,2}\left[\dfrac{\omega_{m,j}}{2}(\hat{p}_{j}^{2}+\hat{q}_{j}^{2})-g_{j}\hat{a}_{\circlearrowright}^{\dagger}\hat{a}_{\circlearrowright}\hat{q}_{j}\right]$
$+\chi\hat{q}_{1}\hat{q}_{2},$
$\hat{H}_{\textrm{dr},\circlearrowright}=$ $i\varepsilon_{d,\circlearrowright}(\hat{a}_{\circlearrowright}^{\dagger}-\hat{a}_{\circlearrowright}),$ (6)
and the associated Born-Markovian master equation is given by Gardiner and Zoller (2000)
$\dot{\rho}_{\circlearrowright}\!=$ $i\!\left[\rho_{\circlearrowright},\hat{H}_{\circlearrowright}\right]\!+\!\frac{\kappa}{2}\mathcal{D}[\hat{a}_{\circlearrowright}]\rho_{\circlearrowright}\!-\!\sum_{j=1,2}\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\rho_{\circlearrowright}\}]\right.$
$\left.+\gamma_{m,j}\bar{n}_{m,j}[\hat{q}_{j},[\hat{q}_{j},\rho_{\circlearrowright}]]\right),$ (7)
where $\hat{a}_{\circlearrowright}$ ($\hat{a}_{\circlearrowright}^{\dagger}$) is the annihilation (creation) operator of the CW mode, and $\rho_{\circlearrowright}$ is the density operator with respect to the CW input case. Note that in the absence of the two-photon parametric driving field, the Bogoliubov transformation has no influence on the system dynamics. Therefore, we apply a normal vacuum reservoir to the CW mode and analyze the dynamical evolution of this system in the original picture.
[…]
For the CCW input case, due to the nonlinear COM interactions in Hamiltonian (3), the system dynamics is difficult to be directly solved. In order to deal with this problem, one can linearize the Hamiltonian by expanding each operator as a sum of its steady-state mean value and a small quantum fluctuation around it under the condition of strong coherent optical driving, i.e., $\hat{a}_{s,\circlearrowleft}=\bar{a}_{s,\circlearrowleft}+\delta\hat{a}_{s,\circlearrowleft}$, $\hat{q}_{j}=\bar{q}_{j}+\delta\hat{q}_{j}$, $\hat{p}_{j}=\bar{p}_{j}+\delta\hat{p}_{j}$. The steady-state mean values of the optical and mechanical modes can be derived by using the master equation (4), i.e.,
$\bar{a}_{s,\circlearrowright}=$ $-\dfrac{(i\Delta_{s}-\dfrac{\kappa}{2})A_{1}+iA_{2}\beta_{p}}{(\Delta_{s}^{2}+\dfrac{\kappa^{2}}{4})-\beta_{p}^{2}}\varepsilon_{d,\circlearrowleft},$
$\bar{q}_{1}=$ $\dfrac{\omega_{m,2}B_{1}-\chi B_{2}}{\omega_{m,1}\omega_{m,2}-\chi^{2}},~{}~{}\bar{p}_{1}=0,$
$\bar{q}_{2}=$ $\dfrac{\omega_{m,1}B_{2}-\chi B_{1}}{\omega_{m,1}\omega_{m,2}-\chi^{2}},~{}~{}\bar{p}_{2}=0,$ (8)
@@
$\alpha_{s}$ $=e^{-i\theta_{d}}\bar{a}_{s,\circlearrowleft}^{\ast 2}+e^{i\theta_{d}}\bar{a}_{s,\circlearrowleft}^{2},$
@@
$B_{1}$ $=\zeta_{s,1}|\bar{a}_{s,\circlearrowleft}|^{2}-\dfrac{\zeta_{p,1}}{2}\alpha_{s}+F_{1},$
$B_{2}$ $=\zeta_{s,2}|\bar{a}_{s,\circlearrowleft}|^{2}-\dfrac{\zeta_{p,2}}{2}\alpha_{s}+F_{2}.$ (9)
Equation (8) indicates that the field amplitude $\bar{a}_{s,\circlearrowleft}$ is not only dependent on the strength of the coherent driving field, but also relies on the squeezing strength and reference angle of the squeezed optical mode. This allows us to regulate the COM coupling strength and break the optical reciprocity of the system by adjusting the squeezing parameters. Moreover, it is also seen that the mean values of the mechanical displacement $\bar{q}_{j}$ are coupled to each other by a factor of $\chi$, which is due to the phonon hopping process between the two mechanical modes.
@@
$\left.-(\Lambda_{j}\hat{a}_{s,\circlearrowleft}^{\dagger}+\Lambda_{j}^{\ast}\hat{a}_{s,\circlearrowleft})\hat{q}_{j}\right]+\chi\hat{q}_{1}\hat{q}_{2},$ (10)
@@
$\dot{\tilde{\rho}}_{\circlearrowleft}\!=$ $i\!\left[\tilde{\rho}_{\circlearrowleft},\tilde{H}_{lin,\circlearrowleft}\right]\!+\!\frac{\kappa}{2}\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}\!-\!\sum_{j=1,2}\!\left(i\dfrac{\gamma_{m,j}}{2}[\hat{q}_{j},\{\hat{p}_{j},\tilde{\rho}_{\circlearrowleft}\}]\right.$
@@
is the effective COM coupling rate, with $G_{j}=g_{j}\bar{a}_{s,\circlearrowleft}$. For notational convenience, we have neglected the symbol ``$\delta"$ in the expression of quantum fluctuation operators in Eqs. (10) and (11). We also emphasize that in the weak COM coupling regime, the steady-state mean value of the optical mode is much larger than those of the mechanical modes, i.e., $|\bar{a}_{s,\circlearrowleft}|\gg|\bar{q}_{j}|$. Under this condition, we have ignored the terms $\hat{a}_{s,\circlearrowleft}^{2}$ and $\hat{a}_{s,\circlearrowleft}^{\dagger 2}$ in Hamiltonian (10), whose coefficient $\beta_{p}$ is dominated by $\bar{q}_{j}$. In the following discussions, for ensuring the validity of Hamiltonian (10), the system parameters have been strictly restricted to satisfy the condition of weak COM coupling.
Figure 2: Squeezing-induced nonreciprocal bipartite and tripartite entanglement. The logarithmic negativity $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$ (a), $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$ (b), $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ (c), and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ (d) for opposite optical input directions are plotted as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$. The left panels of these figures show the bipartite and tripartite entanglement with respect to two specific squeezing reference angles of $\theta_{d}=0.75\pi$ and $\pi$, respectively, with $r_{d}=0.3$. The right panels show in detail the dependence of such entanglement on squeezing reference angle $\theta_{d}$.
@@
Here $\psi=(\hat{X},\hat{Y},\hat{q}_{1},\hat{p}_{1},\hat{q}_{2},\hat{p}_{2})^{T}$ is the vector of optical and mechanical quadrature operators, with its components defined by
$\hat{X}=\dfrac{1}{\sqrt{2}}\left(\hat{a}_{s,\circlearrowleft}^{\dagger}+\hat{a}_{s,\circlearrowleft}\right),~{}~{}~{}\hat{Y}=\dfrac{i}{\sqrt{2}}\left(\hat{a}_{s,\circlearrowleft}^{\dagger}-\hat{a}_{s,\circlearrowleft}\right).$ (14)
Employing the master equation (11), one can obtain the dynamics of arbitrary quantum correlation between the optical and mechanical quadrature operators in the CM $V$. In terms of optical bosonic operator, we define the following quantum correlation functions
$x_{1}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\rangle,~{}x_{2}=\langle\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\rangle,~{}x_{3}=\langle\hat{q}_{1}\hat{q}_{1}\rangle,$
$x_{4}=\langle\hat{p}_{1}\hat{p}_{1}\rangle,~{}x_{5}=\langle\hat{q}_{2}\hat{q}_{2}\rangle,~{}x_{6}=\langle\hat{p}_{2}\hat{p}_{2}\rangle,$
$x_{7}=\langle\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}\rangle,~{}x_{8}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}^{\dagger}\rangle,~{}x_{9}=\langle\hat{q}_{1}\hat{p}_{1}\rangle,$
$x_{10}=\langle\hat{p}_{1}\hat{q}_{1}\rangle,~{}x_{11}=\langle\hat{q}_{2}\hat{p}_{2}\rangle,~{}x_{12}=\langle\hat{p}_{2}\hat{q}_{2}\rangle,$
$x_{13}=\langle\hat{a}_{s,\circlearrowleft}\hat{q}_{1}\rangle,~{}x_{14}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{q}_{1}\rangle,~{}x_{15}=\langle\hat{a}_{s,\circlearrowleft}\hat{p}_{1}\rangle,$
$x_{16}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{p}_{1}\rangle,~{}x_{17}=\langle\hat{a}_{s,\circlearrowleft}\hat{q}_{2}\rangle,~{}x_{18}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{q}_{2}\rangle,$
$x_{19}=\langle\hat{a}_{s,\circlearrowleft}\hat{p}_{2}\rangle,~{}x_{20}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{p}_{2}\rangle,~{}x_{21}=\langle\hat{q}_{1}\hat{q}_{2}\rangle,$
$x_{22}=\langle\hat{p}_{1}\hat{p}_{2}\rangle,~{}x_{23}=\langle\hat{q}_{1}\hat{p}_{2}\rangle,~{}x_{24}=\langle\hat{q}_{2}\hat{p}_{1}\rangle.$ (15)
By grouping together them in a vector, $X=(x_{1},x_{2},\ldots,x_{24})^{T}$, one can obtain its time evolution equation as
$\dfrac{d}{dt}X=M\cdot X+N,$ (16)
@@
$N=$ $\left[0,\kappa,0,2\gamma_{m,1}\bar{n}_{m,1},0,2\gamma_{m,2}\bar{n}_{m,2},0,0,\right.$
$\left.0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\right]^{T},$ (17)
is a vector involving the correlation functions for the input quantum noises. Here the exact expression for the coefficient matrix $M$ and the detailed derivation process of the time evolution equation (16) are too cumbersome, and, for convenience, we have reported them in Appendix C. Notably, by numerically solving Eq. (16) and using the relations between bosonic operators and quadrature operators, one can directly obtain the steady-state CM $V$.
Regarding the verification of bipartite and tripartite entanglement, we adopt the logarithmic negativity $E_{\mathcal{N}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ as the quantitative entanglement measures, respectively, which are defined based on specifying the positivity of the partial transpose of the CM $V$. For continuous-variable (CV) bipartite Gaussian state, the logarithmic negativity $E_{\mathcal{N}}^{\mu|\nu}$ is defined as Adesso et al. (2004)
@@
where $\eta_{0}^{-}\!\equiv\!2^{-1/2}\{\Sigma(V_{\mu\nu})-[\Sigma(V_{\mu\nu})^{2}-4\det\!V_{\mu\nu}]^{1/2}\}^{1/2}$ , with $\Sigma(V_{\mu\nu})\!\equiv\!\det\mathcal{A}_{\mu}+\det\mathcal{B}_{\nu}-2\det\mathcal{C}_{\mu\nu}$, is the minimum symplectic eigenvalue of the partial transpose of a reduced $4\times 4$ CM $V_{\mu\nu}$. The reduced CM $V_{\mu\nu}$ contains the entries of $V$ associated with the selected bipartition $\mu$ and $\nu$, and it can be obtained by removing the rows and columns of the unwanted modes in $V$, whose $2\times 2$ block form is given by
@@
Equation (18) quantifies how much the positivity of the partial transpose condition for separability is violated for the considered Gaussian states, and it is equivalent to Simon’s necessary and sufficient entanglement nonpositive partial transpose criterion (or the related Peres-Horodecki criterion) Simon (2000). The selected bipartition $\mu$ and $\nu$ gets entangled if and only if $\eta_{0}^{-}<1/2$, where $E_{\mathcal{N}}$ has a nonzero value. To further confirm the presence and directionality of Gaussian steering between the two entangled bipartition $\mu$ and $\nu$, we introduce an intuitive and computable quantification of quantum steerability Kogias et al. (2015), which is defined as
$\mathcal{G}^{\mu\rightarrow\nu}=$ $\max\left[0,\dfrac{1}{2}\ln\dfrac{\det\mathcal{A}_{\mu}}{4\det V_{\mu\nu}}\right],$
$\mathcal{G}^{\nu\rightarrow\mu}=$ $\max\left[0,\dfrac{1}{2}\ln\dfrac{\det\mathcal{B}_{\nu}}{4\det V_{\mu\nu}}\right].$ (20)
$\mathcal{G}^{\mu\rightarrow\nu}>0$ ($\mathcal{G}^{\nu\rightarrow\mu}>0$) implies that the bipartite Gaussian state characterized by CM $V_{\mu\nu}$ is steerable from mode $\mu$ ($\nu$) to mode $\nu$ ($\mu$) through Gaussian measurements on mode $\mu$ ($\nu$), where a higher value of $\mathcal{G}$ represents the stronger Gaussian steerability.
The minimum residual contangle, $\mathcal{R}^{\textrm{min}}_{\tau}$, which provides a bona fide quantification of CV tripartite entanglement, is defined as Adesso and Illuminati (2007)

### v2

@@ II Theoretical model and system dynamics
To quantitatively investigate the behavior of quantum entanglement and EPR steering, one should solve the system dynamics by employing the effective mater equation (6). However, due to the nonlinear COM interactions in Hamiltonian (5), the system dynamics is difficult to be directly solved. In order to deal with this problem, one can linearize the Hamiltonian by expanding each operator as a sum of its steady-state mean value and a small quantum fluctuation around it under the condition of strong coherent optical driving, i.e., $\hat{a}_{s,i}=\bar{a}_{s,i}+\delta\hat{a}_{s,i}$, $\hat{q}_{j}=\bar{q}_{j}+\delta\hat{q}_{j}$, $\hat{p}_{j}=\bar{p}_{j}+\delta\hat{p}_{j}$. Since the linearization procedure is identical for the CW and CCW input cases, we here only present the linearization of the effective mater equation (6) regarding the CCW input case. By using the master equation (6), the steady-state mean values of the optical and mechanical modes with respect to the CCW input case are derived as
$\bar{a}_{s,\circlearrowleft}=$ $-\dfrac{(i\Delta_{s}-\dfrac{\kappa}{2})A_{1}+iA_{2}\beta_{p}}{(\Delta_{s}^{2}+\dfrac{\kappa^{2}}{4})-\beta_{p}^{2}}\varepsilon_{d,\circlearrowleft},$
$\bar{q}_{1}=$ $\dfrac{\omega_{m,2}B_{1,\circlearrowleft}-\lambda B_{2,\circlearrowleft}}{\omega_{m,1}\omega_{m,2}-\lambda^{2}},~{}~{}\bar{p}_{1}=0,$
$\bar{q}_{2}=$ $\dfrac{\omega_{m,1}B_{2,\circlearrowleft}-\lambda B_{1,\circlearrowleft}}{\omega_{m,1}\omega_{m,2}-\lambda^{2}},~{}~{}\bar{p}_{2}=0,$ (8)
@@
$\Delta_{s}$ $=\omega_{s}-\beta_{s},$
$\alpha_{s,\circlearrowleft}$ $=e^{-i\theta_{d}}\bar{a}_{s,\circlearrowleft}^{\ast 2}+e^{i\theta_{d}}\bar{a}_{s,\circlearrowleft}^{2},$
@@
$B_{1,\circlearrowleft}$ $=\zeta_{s,1}|\bar{a}_{s,\circlearrowleft}|^{2}-\dfrac{\zeta_{p,1}}{2}\alpha_{s,\circlearrowleft}+F_{1},$
$B_{2,\circlearrowleft}$ $=\zeta_{s,2}|\bar{a}_{s,\circlearrowleft}|^{2}-\dfrac{\zeta_{p,2}}{2}\alpha_{s,\circlearrowleft}+F_{2}.$ (9)
Equation (8) indicates that the field amplitude $\bar{a}_{s,\circlearrowleft}$ is not only dependent on the strength of the coherent driving field, but also relies on the squeezing strength and reference angle of the squeezed optical mode. This allows us to regulate the COM coupling strength in a controllable way by adjusting the squeezing parameters. Moreover, it is also seen that the mean values of the mechanical displacement $\bar{q}_{j}$ are coupled to each other by a factor of $\lambda$, which is due to the phonon hopping process between the two mechanical modes.
@@
$\left.-(\Lambda_{j}\hat{a}_{s,\circlearrowleft}^{\dagger}+\Lambda_{j}^{\ast}\hat{a}_{s,\circlearrowleft})\hat{q}_{j}\right]+\lambda\hat{q}_{1}\hat{q}_{2},$ (10)
@@
$\dot{\tilde{\rho}}_{\circlearrowleft}\!=$ $i\!\left[\tilde{\rho}_{\circlearrowleft},\tilde{H}_{lin,\circlearrowleft}\right]$
$+\!\frac{\kappa}{2}(N_{s,\circlearrowleft}+1)\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}\!+\!\frac{\kappa}{2}N_{s,\circlearrowleft}\mathcal{D}[\hat{a}_{s,\circlearrowleft}^{\dagger}]\tilde{\rho}_{\circlearrowleft}$
$-\frac{\kappa}{2}M_{s,\circlearrowleft}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}-\frac{\kappa}{2}M_{s,\circlearrowleft}^{\ast}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}$
@@
is the effective COM coupling rate, with $G_{j}=g_{j}\bar{a}_{s,\circlearrowleft}$. By defining an enhancement factor $\Pi_{j}=|\Lambda_{j}/G_{j}|$, it is seen that with proper squeezing parameters, the COM coupling rate can be effectively enhanced [see Fig. 1(b)]. Note that for notational convenience, we have neglected the symbol ``$\delta"$ in the expression of quantum fluctuation operators in Eqs. (10) and (11). We also emphasize that in the weak COM coupling regime, the steady-state mean value of the optical mode is much larger than those of the mechanical modes, i.e., $|\bar{a}_{s,\circlearrowleft}|\gg|\bar{q}_{j}|$. Under this condition, we have ignored the terms $\hat{a}_{s,\circlearrowleft}^{2}$ and $\hat{a}_{s,\circlearrowleft}^{\dagger 2}$ in Hamiltonian (10), whose coefficient $\beta_{p}$ is dominated by $\bar{q}_{j}$. In the following discussions, for ensuring the validity of Hamiltonian (10), the system parameters have been strictly restricted to satisfy the condition of weak COM coupling.
Figure 2: Phase-selective generation of optomechanical entanglement. (a),(c) Plot of the logarithmic negativity $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{1}}}$ and $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{2}}}$ versus the scaled optical detuning $\Delta_{c}/\omega_{m}$ for opposite optical input directions with $r_{d}=0.1$ and $\theta_{d}=\pi$. (b),(d) Density plot of $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ as functions of $\delta r_{\circlearrowright}$ and $\delta\theta_{\circlearrowright}$, with $\delta r_{\circlearrowleft}=0$ and $\delta\theta_{\circlearrowleft}=\pi$. (e),(f) Plot of the logarithmic negativity $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{1}}}$ and $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{2}}}$ versus the mechanical environment temperature $T$ for opposite optical input directions with $r_{d}=0.1$, $\theta_{d}=\pi$, and $\Delta_{c}=1$.
@@
Here $\psi=(\hat{X}_{s,\circlearrowleft},\hat{Y}_{s,\circlearrowleft},\hat{q}_{1},\hat{p}_{1},\hat{q}_{2},\hat{p}_{2})^{T}$ is the vector of optical and mechanical quadrature operators, with its components defined by
$\hat{X}_{s,\circlearrowleft}=\dfrac{1}{\sqrt{2}}\left(\hat{a}_{s,\circlearrowleft}^{\dagger}+\hat{a}_{s,\circlearrowleft}\right),~{}\hat{Y}_{s,\circlearrowleft}=\dfrac{i}{\sqrt{2}}\left(\hat{a}_{s,\circlearrowleft}^{\dagger}-\hat{a}_{s,\circlearrowleft}\right).$ (14)
Employing the linearized master equation (11), one can obtain the dynamics of arbitrary quantum correlation between the optical and mechanical quadrature operators in the CM $V$. In terms of optical bosonic operator, we define the following quantum correlation functions
$x_{1}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}\rangle,~{}~{}x_{2}=\langle\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}^{\dagger}\rangle,~{}~{}x_{3}=\langle\hat{q}_{1}\hat{q}_{1}\rangle,$
$x_{4}=\langle\hat{p}_{1}\hat{p}_{1}\rangle,~{}~{}x_{5}=\langle\hat{q}_{2}\hat{q}_{2}\rangle,~{}~{}x_{6}=\langle\hat{p}_{2}\hat{p}_{2}\rangle,$
$x_{7}=\langle\hat{a}_{s,\circlearrowleft}\hat{a}_{s,\circlearrowleft}\rangle,~{}~{}x_{8}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{a}_{s,\circlearrowleft}^{\dagger}\rangle,~{}~{}x_{9}=\langle\hat{q}_{1}\hat{p}_{1}\rangle,$
$x_{10}=\langle\hat{p}_{1}\hat{q}_{1}\rangle,~{}~{}x_{11}=\langle\hat{q}_{2}\hat{p}_{2}\rangle,~{}~{}x_{12}=\langle\hat{p}_{2}\hat{q}_{2}\rangle,$
$x_{13}=\langle\hat{a}_{s,\circlearrowleft}\hat{q}_{1}\rangle,~{}~{}x_{14}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{q}_{1}\rangle,~{}~{}x_{15}=\langle\hat{a}_{s,\circlearrowleft}\hat{p}_{1}\rangle,$
$x_{16}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{p}_{1}\rangle,~{}~{}x_{17}=\langle\hat{a}_{s,\circlearrowleft}\hat{q}_{2}\rangle,~{}~{}x_{18}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{q}_{2}\rangle,$
$x_{19}=\langle\hat{a}_{s,\circlearrowleft}\hat{p}_{2}\rangle,~{}~{}x_{20}=\langle\hat{a}_{s,\circlearrowleft}^{\dagger}\hat{p}_{2}\rangle,~{}~{}x_{21}=\langle\hat{q}_{1}\hat{q}_{2}\rangle,$
$x_{22}=\langle\hat{p}_{1}\hat{p}_{2}\rangle,~{}~{}x_{23}=\langle\hat{q}_{1}\hat{p}_{2}\rangle,~{}~{}x_{24}=\langle\hat{q}_{2}\hat{p}_{1}\rangle.$ (15)
By grouping together these quantum correlations in a vector $X=(x_{1},x_{2},\ldots,x_{24})^{T}$ and employing the linearized effective master equation (11), one can obtain its time evolution equation as
$\dfrac{d}{dt}X=A\cdot X+b,$ (16)
@@
$b=$ $\left[\kappa N_{s},\kappa(N_{s}+1),0,2\gamma_{m,1}\bar{n}_{m,1},0,2\gamma_{m,2}\bar{n}_{m,2},\kappa M_{s}^{\ast},\right.$
$\left.\kappa M_{s},0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\right]^{T},$ (17)
is a vector involving the correlation functions for the input quantum noises. Here the exact expression for the coefficient matrix $A$ and the detailed derivation process of the time evolution equation (16) are too cumbersome, and, for convenience, we have reported them in Appendix C. Notably, by numerically solving Eq. (16) and using the relations between bosonic operators and quadrature operators, one can directly obtain the steady-state CM $V$.
Regarding the verification of bipartite and tripartite entanglement, we adopt the logarithmic negativity $E_{\mathcal{N}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ as the quantitative entanglement measures, respectively, which are defined based on specifying the positivity of the partial transpose of the CM $V$. For continuous-variable (CV) bipartite Gaussian state, the logarithmic negativity $E_{\mathcal{N}}^{\mu|\nu}$ is defined as [15]
@@
where $\eta_{0}^{-}\!\equiv\!2^{-1/2}\{\Sigma(V_{\mu\nu})-[\Sigma(V_{\mu\nu})^{2}-4\det\!V_{\mu\nu}]^{1/2}\}^{1/2}$, with $\Sigma(V_{\mu\nu})\!\equiv\!\det\mathcal{A}_{\mu}+\det\mathcal{B}_{\nu}-2\det\mathcal{C}_{\mu\nu}$, is the minimum symplectic eigenvalue of the partial transpose of a reduced $4\times 4$ CM $V_{\mu\nu}$, with $\mu$ and $\nu$ describing the selected modes under consideration ($\mu,\,\nu=\hat{a}_{s,\circlearrowright},\hat{a}_{s,\circlearrowleft},\hat{q}_{1},\hat{q}_{2}$). The reduced CM $V_{\mu\nu}$ contains the entries of $V$, and it can be obtained by selecting the rows and columns associated with modes $\mu$ and $\nu$ from $V$, whose $2\times 2$ block form is given by
@@
Equation (18) quantifies how much the positivity of the partial transpose condition for separability is violated for the considered Gaussian states, which is equivalent to Simon’s necessary and sufficient entanglement nonpositive partial transpose criterion (or the related Peres-Horodecki criterion) [14]. Note that the selected modes $\mu$ and $\nu$ gets entangled if and only if $\eta_{0}^{-}<1/2$, where $E_{\mathcal{N}}$ has a nonzero value.
Figure 3: Phase-selective generation of asymmetric optomechanical EPR steering. The quantum steerability $\mathcal{S}^{\hat{a}_{s,i}\rightarrow\hat{q}_{1}}$ (a) and $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{a}_{s,i}}$ (b) for opposite optical input directions are plotted as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing parameter $r_{d}$, with $\delta r_{\circlearrowleft}=\textrm{0}$, $\delta r_{\circlearrowright}=\textrm{0.4}$, and $\delta\theta_{\circlearrowleft}=\delta\theta_{\circlearrowright}=\pi$. With the increase of $r_{d}$, EPR steering can be driven from no-way regime to one-way regime for the CCW input case but not for the CW input case, implying that the asymmetry of EPR steering are effectively controlled by the phase-matched condition.
The minimum residual contangle, $\mathcal{R}^{\textrm{min}}_{\tau}$, which provides a bona fide quantification of CV tripartite entanglement, is defined as [16]